PROBING THE ELECTRONIC STRUCTURE OF YbCl USING LASER SPECTROSCOPY

\underline{TODD C. MELVILLE} and JOHN A. COXON, Department of Chemistry, Dalhousie University, Halifax, N.S. Canada B3H 4J3.

The high resolution spectrum of the A^{2}Π-X^{2}Σ^{+} transition of YbCl has been recorded near 550 nm using laser excitation spectroscopy. The output of a Coherent 699-29 ring dye laser operating in single frequency mode was used to obtain Doppler-limited spectra. Selective detection of laser induced fluorescence was utilized to record spectra with an accuracy of 0.003 cm^{−1}. Unequivocal assignment of the rotational numbering was obtained using resolved fluorescence spectra. Data from two isotopomers, ^{172}Yb^{35}Cl and ^{174}Yb^{35}Cl, have been employed in a least-squares fit of sets of molecular constants. The rotational analysis of the A - X system will be discussed.